### v1

Abstract
Quadcopters have been studied for decades thanks to their maneuverability and capability of operating in a variety of circumstances. However, quadcopters suffer from dynamical nonlinearity, actuator saturation, as well as sensor noise that make it challenging and time consuming to obtain accurate dynamic models and achieve satisfactory control performance. Fortunately, deep reinforcement learning came and has shown significant potential in system modelling and control of autonomous multirotor aerial vehicles, with recent advancements in deployment, performance enhancement, and generalization. In this paper, an end-to-end deep reinforcement learning-based controller for quadcopters is proposed that is secure for real-world implementation, data-efficient, and free of human gain adjustments. First, a novel actor-critic-based architecture is designed to map the robot states directly to the motor outputs. Then, a quadcopter dynamics-based simulator was devised to facilitate the training of the controller policy. Finally, the trained policy is deployed on a real Crazyflie nano quadrotor platform, without any additional fine-tuning process. Experimental results show that the quadcopter exhibits satisfactory performance as it tracks a given complicated trajectory, which demonstrates the effectiveness and feasibility of the proposed method and signifies its capability in filling the simulation-to-reality gap.
keywords:
deep reinforcement learning, quadcopter controller, end-to-end, sim2real, and actor-critic network.
1 Introduction
Unmanned aerial vehicles (UAVs), in particular quadrotors, are increasingly prevalent research subjects due to their flexibility and autonomy [1, 2, 3]. With hovering capabilities and vertical take-off and landing (VTOL), a plethora of applications including search and rescue, infrastructure assessment, and urban air mobility [4, 5]. Precision and agility in flying movements are essential in these applications. The dynamics of quadrotors are highly nonlinear and often hard to model specific system, which presents a considerable challenge in terms of stabilization control. Conventional cascaded hierarchy controls require domain expertise and engineering to be adjusted for changing hardware and utilization circumstances. Moreover, this method often takes extensive setup and experiment time, along with parameter optimization [6, 7].
Recent developments in machine learning and deep learning have proven efficient in addressing various complicated problems [8, 9]. Robot control is a decision-making problem that can be characterized as a Markov Decision Process (MDP), in contrast to supervised learning, in which labels tend to be not directly exist. In order to deal with MDPs, reinforcement learning has been utilized to train policies for complicated end-to-end control problems in simulation [10, 11, 12]. Additionally, they demonstrate encouraging results in learning tasks involving continuous state/action space [10], which have a close relationship to quadrotor control [13]. Compared with typical optimization methods, this technique eliminates the necessity for a specified controller structure, which might constrain agent performance and increase human effort. Despite results achieved in simulation are incredible, the reality gap between simulation and real-world systems is a notable challenge to the deployment of trained control policies. This is essentially due to model imperfections, inaccurate state observations, uncertainty from observation and action, and other disturbances [14, 15].
In this research, an end-to-end reinforcement learning based quadcopter controller is proposed to deal with the aforementioned problems. The aim is to create a controller that is extremely data efficient, free of human gain adjustments, and is safe for real-world implementation. First, a novel actor-critic-based architecture is designed to directly translate quadcopter states to motor Revolutions Per Minute (RPM) outputs without any supplementary pre-configured controllers. Then, a python simulated environment was created using OpenAI Gym [16], specifically designed for training transferable policies. The setting is based on a lightweight quadcopter platform known as Crazyflie 2.1111Crazyflie 2.1: https://www.bitcraze.io/products/crazyflie-2-1/ [17], which has been equipped with thrust upgraded motors and propellers. Finally, practical flying tests were conducted to validate the adaptation of the sim2real strategy. The preliminary results indicate that using a neural network, trained with reinforcement learning methods in a simulated environment, is both effective and feasible for comprehensive control of the drone.
The remainder of this article are organized in the following manner. Section 2  provides an overview of the methodologies including the quadcopter dynamics simulated environment, policy networks, and the training strategy. The experiments, involving the system’s configuration and the obtained results, are shown in section 3  Furthermore, in section 4  , paper concludes by outlining potential future research and development opportunities.
2 Methodology
The diagram of end-to-end deep reinforcement learning quadcopter controller algorithm is demonstrated in Fig. 1. The quadcopter dynamics simulator is first established, followed by the designing of an actor-critic-based policy network that directly maps the states to the RPM commands of the motor. Quaternions are generally adopted to demonstrate orientation since they provide a global and compact expression. However, there is a specific drawback in our case, which is the existence of two figures that indicates the same rotation (i.e. $q=-q$). Therefore, we must either increase the amount of training data by two-fold or accept the issue of having a discontinuous function when we limit our domain to one hemisphere of $S^{3}$. The rotation matrix is a representation that is highly redundant, but it is also simple and free from such issues. Hence, in order to eliminate the ambiguity that rose from the quaternion’s double coverage of the space of rotations, we convert them to a nine-element rotation matrix $R$.
Moreover, we observe that the delay in the step-response is significant and is an effect of the low-pass behavior indicated by the motors. This behavior greatly affects the dynamics of the Crazyflie nano quadrotor used in this study. According to the manufacturer’s measurements [^2] Crazyflie motor step response: https://www.bitcraze.io/2015/02/measuring-propeller-rpm-part-3/, we have determined through empirical analysis that a time interval of $0.15$ seconds would be sufficient for sim2real transference. These delays are considerably longer than the typical control interval of low-level controllers in Crazyflie. Therefore, they result in actions that affect the state only after $5$ to $25$ control steps. In order to tackle the major amount of partial observability, we include the history of control actions into the observation. This is a proprioceptive measurement that can be easily done in software without the need of additional hardware. In addition, noise was introduced into the observation to imitate imperfections in the sensor feedback of the reality platform. The unexpected noise values have a standard deviation of $0.001$ for position and orientation, and $0.002$ for linear velocity and angular velocity, respectively.
The observations are $18+4\cdot N_{H_{a}}$ dimensional, where $N_{H_{a}}$ is the number of action history, described as follows:
$s_{t}=[p_{t},R_{t},v_{t},\omega_{t},H_{a}]$ (1)
where $p_{t}$ denotes the position; $R_{t}$ is the rotation matrix; $v_{t}$ is linear velocity; $\omega_{t}$ is the angular velocity; $H_{a}$ is the history of actions.
The actions comprise the RPM setpoints for each motor inside of the continuous space range of $[\mathrm{min}_{a},\mathrm{max}_{a}]$ shown as below:
$a_{t}=[r_{1},r_{2},r_{3},r_{4}]=\pi_{\Gamma}(s_{t})$ (2)
Figure 1: Actor-critic-based end-to-end deep reinforcement learning quadcopter controller approach diagram.
2.1 Quadcopter dynamics simulator
We create a simulator that relies on the quadcopter dynamics model to train the policy network. In previous studies, a number of methods have been proposed and demonstrated to determine the dynamics model of the quadcopter using simulations and experiments [7, 6]. Let $\Xi=[\phi,\theta,\psi]^{T}\in\mathbb{R}$3 denotes the quadcopter’s attitude including roll, pitch, and yaw angles in the Earth frame {E} as shown in Fig. 2. The vehicle’s angular velocity in the body frame {B} denoted as $\omega_{b}\in\mathbb{R}$3. The position and velocity of the quadcopter along $\mathrm{x}$, $\mathrm{y}$, $\mathrm{z}$-axis of {E} are indicated by $p=[x,y,z]^{T}\in\mathbb{R}$3, and $v=\dot{p}$, respectively. $J=\mathrm{diag}(I_{xx},I_{yy},I_{zz})$ $\in\mathbb{R}$3×3 denotes the moment of inertia in {B}. The quadcopter’s dynamics model can be described by the following equations:
$\omega_{b}$ $=\mathrm{R}(\Xi)\dot{\Xi}$ (3)
$m\dot{v}$ $=mg+\mathrm{R}(\Xi)u_{3}F_{\sum}$ (4)
$J\dot{\omega}_{b}$ $=(J\omega_{b})\times\omega_{b}+\tau$ (5)
where $m$ denotes the total mass of the quadcopter; $g\in\mathbb{R}$3 is the gravitational acceleration vector; $\mathrm{u}_{3}=[0,0,1]^{T}$; $F_{\sum}$ is the total thrust force; $R(\Xi)$ is the rotation matrix from {B} to {E}, which is defined as:
$R(\Xi)=\begin{bmatrix}1&0&\mathrm{-sin}(\theta)\\ 0&\mathrm{cos}(\theta)&\mathrm{sin}(\phi)\mathrm{cos}(\theta)\\ 0&\mathrm{-sin}(\theta)&\mathrm{cos}(\phi)\mathrm{cos}(\theta)\end{bmatrix}$ (6)
Figure 2: The configuration of the quadcopter and coordinate systems with a body frame {B} (x, y, z) and the Earth frame {E} (N, E, D).
The control torque $\tau=[\tau_{1},\tau_{2},\tau_{3}]^{T}$ is generated by four propellers attached to four motors as follows:
$\begin{cases}\tau_{1}=LC_{b}(\Omega_{2}^{2}-\Omega_{4}^{2})\cr\tau_{2}=LC_{b}(\Omega_{3}^{2}-\Omega_{1}^{2})\cr\tau_{3}=C_{d}C_{b}(-\Omega_{1}^{2}+\Omega_{2}^{2}-\Omega_{3}^{2}+\Omega_{4}^{2})\end{cases}$ (7)
where $L$ represents the arm length of the quadcopter; ${\Omega}_{i}$ denotes the $i$-th motor’s rotary speed $(i=1,2,3,4)$; $C_{b}$ and $C_{d}$ consequently represent thrust and drag coefficient.
2.2 Networks architecture
We take advantage of the Twin Delayed Deep Deterministic policy gradient (TD3) [18], an off-policy reinforcement learning approach that provides enhanced sample complexity.
There are two networks, specifically a value network and a policy network, adopted for training. Both networks take the state as an input. We involve a fully-connected neural network to represent a policy. The policy network is formed by two hidden layers, each containing $64$ neurons. The activation function employed in these layers is $\mathrm{tanh}$, whereas the output layer uses linear activation. The value network has a same structure and is trained for predicting the state-action value function.
2.3 Rewards
For the initial state distribution, we sample uniformly from the following sets: The position is sampled inside a $0.2\mathrm{m}$ box centered around the origin location. The orientation is represented by the rotation matrix $R$, which belongs to the entire $\mathrm{SO}(3)$ group and has a maximum angle of $90$ degrees. The linear velocity has a maximum magnitude of $1$ $\mathrm{m/s}$, while the angular velocity has a maximum value of $1$ $\mathrm{rad/s}$. In order to solve the issue of “learning to terminate” [19], we use a negative squared reward function and includes an additional constant to encourage survival as follows:
$r(s_{t},a_{t},s_{t+1})=$ $\lambda_{s}-\eta_{p}\|p_{t}\|^{2}_{2}-\eta_{R}(1-R_{t}^{2})$
$-\eta_{v}\|v_{t}\|^{2}_{2}-\delta_{a}\|a_{t}-\delta_{ab}\|^{2}_{2}$ (8)
where $\lambda_{s}$ is the survival bonus; $\eta_{p}$, $\eta_{R}$, $\eta_{v}$, and $\delta_{a}$ are the position, orientation, linear velocity, and action weights, respectively; $\delta_{ab}$ is the action baseline weight.
The survival bonus is configured with a high value in order to encourage the agent to maintain its life. The agent is penalized for deviating from the target state based on its position, orientation, linear velocity, and action terms. The position and orientation components have the highest cost coefficients due to their significant impact on the overall outcome. The other cost terms are chosen minimized to prevent the agent from executing unreasonable actions, which might result in crashing or straying from the desired state.
3 Experiments
3.1 Systems Configurations
Table 1: QUADCOPTER PARAMETERS USED FOR THE SIMULATOR.
Parameter Value Unit
$m$ 0.033 kg
$g$ $[0,0,-9.81]^{T}$ m/s2
$L$ 0.028 m
$C_{b}$ 0.0059 Ns2
$C_{d}$ 9.18 ×$\times 10^{-7}$ kg/rad
$I_{xx}$ 16.57 ×$\times 10^{-6}$ kgm2
$I_{yy}$ 16.66 ×$\times 10^{-6}$ kgm2
$I_{zz}$ 29.26 ×$\times 10^{-6}$ kgm2
[…]
The experiments were conducted in a $5\times 4\times 2.5\mathrm{m}^{3}$ flying space at the Guidance, Navigation and Control Lab (GNC Lab) at Sejong University with the VICON Motion Capturing System consisting of $12$ cameras as illustrated in Fig. 3a). It is assumed that we have access to actually precise estimates of the quadrotor’s position, orientation, linear velocity, and angular velocity. The VICON system supplies the position and linear velocity data to a base station computer at a frequency of $100\mathrm{Hz}$, whereas the updates for orientation angle state estimate occur at $1\mathrm{kHz}$. The state estimate is obtained with an extended Kalman filter (EKF) [21] that combines data from the on-board IMU with motion capture information.
The control policy is a C-function that is automatically generated from the trained neural network model in PyTorch [^4] https://pytorch.org/. Next, we included the functionality to output the RPMs of all four motors in the Crazyflie firmware. The Crazyswarm API [22] is used to handle communication with the drone. The desired positions are sent to the quadrotor at a rate of $50\mathrm{Hz}$ via a $2.4\mathrm{GHz}$ radio dongle.
Figure 3: Experimental conditions. a) VICON positioning system, b) Crazyflie 2.1 with thrust upgrade kit
3.2 Experimental results
[…]
Table 2: PARAMETERS OF REWARDS FUNCTION.
Parameter Value Description
$\lambda_{s}$ 2 Survival bonus
$\eta_{p}$ 2.5 Position weight
$\eta_{R}$ 2.5 Orientation weight
$\eta_{v}$ 0.05 Linear velocity weight
$\delta_{a}$ 0.05 Action weight
$\delta_{ab}$ 0.35 Action baseline
Figure 4: Real flight tracking performance of circle trajectory with a radius of $1\mathrm{m}$ in $T$=$6$ seconds.
Figure 5: Real flight tracking of a circle trajectory in the $\mathrm{x}$, $\mathrm{y}$, and $\mathrm{z}$ axes. Left: position tracking performance; Right: linear velocity tracking performance.
We evaluate the performance of our policy on the real Crazyflie quadrotor to follow a circle trajectory start at $\mathrm{x}$ = $1$ and $\mathrm{y}$ = 00 with a radius of $1\mathrm{m}$ in $T$=$6$ seconds. The formula of the circle trajectory is defined as:
$p^{d}(t)=\begin{bmatrix}x^{d}_{t}\\ y^{d}_{t}\\ z^{d}_{t}\\ \end{bmatrix}=\begin{bmatrix}1\\ 0\\ 1\\ \end{bmatrix}=\begin{bmatrix}cos(2\pi t/T)\\ sin(2\pi t/T)\\ 0\\ \end{bmatrix}$ (9)
where $T$ is the cycle time.
The result of tracking performance is demonstrated in Fig. 4. There is a minor tracking error but it is not a significant amount. The trained policy is able to track circle trajectories in $6$ seconds as in Fig. 5 where it reaches up to $1.8\mathrm{m/s}$. The real world experiments video can be found at: http://bit.ly/gnc_drl_quad_controller.
We compared our proposed approach with different traditional controllers, including Proportional-Integral-Derivative (PID) and Mellinger [24], utilizing same trajectory tracking task, hardware configuration, and Crazyswarm default gains. Compared to the other controllers our trained controller directly generates RPMs as output and hence does not take advantage battery voltage compensation, so we consider both the Root-Mean-Square Error (RMSE) with and without the $\mathrm{z}$ component as $\bar{e}$ and $\bar{e}_{xy}$ shown in Table 3, which are calculated as:
$\bar{e}$ $=\sqrt{\frac{1}{3}\left((x-x^{d})^{2}+(y-y^{d})^{2}+(z-z^{d})^{2}\right)}$ (10)
$\bar{e}_{xy}$ $=\sqrt{\frac{1}{2}\left((x-x^{d})^{2}+(y-y^{d})^{2}\right)}$ (11)
where $x^{d}$, $y^{d}$, $z^{d}$ are the desired position, and $x$, $y$, $z$ are the actual position.
Our controller policy has the capability of taking-off from the ground, which means overcoming the impact of near-ground airflow effects. While tracking the given trajectory, the Mellinger controller outperforms the PID controller due to the inclusion of an integral component, which aids in reducing steady state error.
Although our policy does not have integral part memory, but it is equivalent to the Mellinger controller in terms of distance with $0.18\mathrm{m}$ and has a smaller error in the $\mathrm{xy}$-plane with $0.16\mathrm{m}$.
Table 3: Comparison of tracking error when tracking the circle trajectory of different controllers.
Controller Tracking error (m)
$\bar{e}$ $\bar{e}_{xy}$
PID 0.68 0.67
Mellinger [24] 0.18 0.18
Ours 0.18 0.16
4 Conclusion
This research presents a reinforcement learning scheme for a quadrotor controller that generates RPMs directly based on the vehicle’s states. Once trained with the simulator on a computer, the policy is directly utilized on real-world Crazyflie platforms for carrying out real flying tests in the VICON positioning system, without necessity for expert parameter adjustments. The presented technique has been evaluated by experimental results, which indicate its efficiency and practicality. The trained policy is able of operating onboard and demonstrates precise trajectory tracking capabilities.
There are still potential to further development in order to achieve even more outstanding results. In future works, it is essential to enhance the robustness by extending the policy to take into account the changes in system parameters or environmental disturbances, such as variations in battery levels or wind gusts.
ACKNOWLEDGEMENT
This work was supported by Future Space Navigation & Satellite Research Center through the National Research Foundation funded by the Ministry of Science and ICT, the Republic of Korea (2022M1A3C2074404).

### v2

@@
 Hence, in order to eliminate the ambiguity that rose from the quaternion’s double coverage of the space of rotations, we convert them to a nine-element rotation matrix $R$.
-Moreover, we observe that the delay in the step-response is significant and is an effect of the low-pass behavior indicated by the motors.
+Moreover, it is observed that the delay in the step-response is significant and is an effect of the low-pass behavior indicated by the motors.
 This behavior greatly affects the dynamics of the Crazyflie nano quadrotor used in this study.
@@
 These delays are considerably longer than the typical control interval of low-level controllers in Crazyflie.
-Therefore, they result in actions that affect the state only after $5$ to $25$ control steps.
+Therefore, they result in actions that affect the state only after $5$ to $25$ control steps [18].
 In order to tackle the major amount of partial observability, we include the history of control actions into the observation.
 This is a proprioceptive measurement that can be easily done in software without the need of additional hardware.
-In addition, noise was introduced into the observation to imitate imperfections in the sensor feedback of the reality platform.
+In addition, noise was introduced into the observation to imitate imperfections in the sensor feedback of the reality platform [18].
 The unexpected noise values have a standard deviation of $0.001$ for position and orientation, and $0.002$ for linear velocity and angular velocity, respectively.
-The observations are $18+4\cdot N_{H_{a}}$ dimensional, where $N_{H_{a}}$ is the number of action history, described as follows:
+The observations are $18+4\cdot N_{H_{a}}$ dimensional, where $N_{H_{a}}$ is the number of action history [18], described as follows:
 $s_{t}=[p_{t},R_{t},v_{t},\omega_{t},H_{a}]$ (1)
@@
 In future works, it is essential to enhance the robustness by extending the policy to take into account the changes in system parameters or environmental disturbances, such as variations in battery levels or wind gusts.
-ACKNOWLEDGEMENT
-This work was supported by Future Space Navigation & Satellite Research Center through the National Research Foundation funded by the Ministry of Science and ICT, the Republic of Korea (2022M1A3C2074404).